# A Mixed-Order Phase Transition in Continuous-Variable Quantum Networks

### v1

@@ B G-G entanglement concentration
Actually, Lemma 1 provides a sufficient condition to convert two TMSVSs $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ into another one with squeezing parameter $r_{1}^{\prime}>r_{1}$ for the case of $r_{2}^{\prime}=0$. To be exact, the maximum value $r$ of $r_{1}^{\prime}$ that satisfies Eq. (I.9) is given by $r=\sinh^{\mbox{-}1}(\sinh{r_{1}}\cosh{r_{2}})>r_{1}$ and $r_{2}^{\prime}=0$, implying that entanglement is concentrated. Thus, consider a $(2+2)$-mode pure Gaussian state $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ with decreasingly ordered squeezing vectors $\vec{r}^{\downarrow}=(r_{1},r_{2})$. This state can be deterministically converted into a single TMSVS $|\psi^{r}\rangle$ by non-Gaussian LOCC, where
$\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}.$ (I.10)
Equation (I.10) is the building block for scaling the concentration protocol, as demonstrated in Table. 2. Let $|\Psi\rangle$ be a pure Gaussian state and, w.l.o.g., write $|\Psi\rangle=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$ $(r_{1}\geqslant r_{2}\geqslant\cdots\geqslant r_{K}>0)$ with squeezing vector $\vec{r}^{\downarrow}=\left(r_{1},r_{2},\dots,r_{K}\right)$ by Ref. [64]. By iteratively applying Eq. (I.10) $K-1$ times on $|\Psi\rangle$ (Fig. S2), the pure Gaussian state $|\Psi\rangle$ is deterministically concentrated into a single TMSVS. Only LOCC is used throughout the entire process, giving rise to the following theorem:
@@
$\sinh{r}=\sinh{r_{1}}\prod_{k=2}^{K}\cosh{r_{k}}.$ (I.11)
Proof.
@@
Figure S2: Continuous-variable entanglement concentration. Consider $K$ parallel TMSVSs written as the tensor product: $\bigotimes\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$. Step $a_{1}$: The TMSVS $|\psi^{r_{1}^{\prime}}\rangle$ is obtained by executing the TMSVS entanglement concentration protocol on $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$; Step $a_{2}$: Performing the scheme again on $|\psi^{r_{1}^{\prime}}\rangle\otimes|\psi^{r_{3}}\rangle$ yields the TMSVS $|\psi^{r_{2}^{\prime}}\rangle$; and so on. This eventually results in a single TMSVS $|\psi^{r}\rangle$, where the squeezing parameter $r$ satisfies Eq. (I.11).
Optimal concentration order.—In general, depending on which two modes are first concentrated, the prefactor in Eq. (I.11) could be $\sinh{r_{k}}$ for any $k$ other than $\sinh{r_{1}}$. However, the optimal entanglement of $|\psi^{r}\rangle$ is obtained by first concentrating the mode of the highest entanglement (largest squeezing parameter, $r_{1}$):
@@
The squeezing parameter $r$ of the final TMSVS $|\psi^{r}\rangle$ is maximized if and only if the TMSVS with the largest squeezing parameter ($\max\limits_{1\leqslant k\leqslant K}r_{k}$) is among the first two TMSVSs to be concentrated. The corresponding optimal $r$ is given by $\sinh r=(\tanh{\max\limits_{1\leqslant k\leqslant K}r_{k})}\left(\prod_{k=1}^{K}\cosh r_{k}\right)$.
Proof.
@@
where $r^{\prime}=\text{arcsinh}(\sinh r_{2}\cosh r_{3})$. Since $r_{1}\geqslant r_{k}$ for $k=2,3$, we have $\sin r_{1}\cosh r_{2}\cosh r_{3}\geqslant\cosh r_{1}\sinh r^{\prime}$. Furthermore, $\sin r_{1}\cosh r_{2}\cosh r_{3}\geqslant\sinh r_{1}\cosh r^{\prime}$ holds, due to
@@
Case 2 ($K>3$). Assume that this theorem holds true for $K=n\ (n\geqslant 3)$. Then, for $K=n+1$, $|\psi^{r_{1}}\rangle$ must either be the first to concentrate (since it is the largest among the first $n$ TMSVSs to be concentrated), or the last to concentrate. In the latter case, $|\psi^{r_{2}}\rangle$ (with the second largest $r_{2}$), which is now the largest among the first $n$ TMSVSs, must be the first to be concentrated. This gives rise to
@@
where $r^{\prime\prime}=\text{arcsinh}\left(\sinh r_{2}\prod_{k=3}^{n+1}\cosh r_{k}\right)$. Because $r_{1}\geqslant r_{2}$ and
@@
$\leqslant$ $\prod_{k=2}^{n+1}\cosh r_{k},$
we have $\sinh r_{1}\prod_{k=2}^{n+1}\cosh r_{k}\geqslant\cosh r_{1}\sinh r^{\prime\prime}$ and $\sinh r_{1}\prod_{k=2}^{n+1}\cosh r_{k}\geqslant\sinh r_{1}\cosh r^{\prime\prime}$, respectively. It is therefore optimal to concentrate $|\psi^{r_{1}}\rangle$ first. ∎
To express Eq. (I.11) in terms of entanglement measure, let $\chi_{k}$ and $\chi$ be the ratio negativity of input TMSVS $|\psi^{r_{k}}\rangle$ and the output TMSVS $|\psi^{r}\rangle$, respectively. We have:
$\frac{\chi^{2}}{1-\chi^{2}}=\frac{\max\limits_{1\leqslant k\leqslant K}\chi_{k}^{2}}{\prod^{K}_{k=1}\left(1-\chi_{k}^{2}\right)},$ (I.12)
@@
$\text{para}(\chi_{1},\chi_{2},...):=\chi=\frac{\max\limits_{1\leqslant k\leqslant K}\chi_{k}}{\sqrt{\max\limits_{1\leqslant k\leqslant K}\chi_{k}^{2}+\prod_{k=1}^{K}(1-\chi_{k}^{2})}}.$
Section II Negativity Percolation in Non-Series-Parallel QNs
For non-series-parallel QNs, the negativity percolation theory (NegPT) based on the ratio negativity is defined using approximate star-mesh transforms (more details are provided in Ref. [28]). A star-mesh transform [93] can establish a local equivalence of the connectivity between an $(s+1)$-node star graph and an $s$-node complete graph, based only on series and parallel rules. Note that the star-mesh transform is generally irreversible. Specifically, while a star graph with arbitrary weights can most likely be transformed into a complete graph with matching weights, the inverse transformation from a complete graph with arbitrary weights back to a star graph is usually not possible. An exception to this is when $s=3$, in which case the star-mesh transform is reversible and is known as the “Y-$\Delta$/$\Delta$-Y” transforms.
@@
Figure S3: Wheatstone bridge. S3(a) The original QN. S3(b) The QN after the “Y-$\Delta$” transform. S3(c) The final QN.
A Wheatstone bridge
From Eqs. (I.5) and (B), we have
@@ A Wheatstone bridge
The simplest non-series-parallel network is a 4-node network: a Wheatstone bridge, as shown in Fig. S3(a). Let each link in the QN be a TMSVS with identical ratio negativity $\chi$. We apply the star-mesh transform to $R_{2}$, namely the “Y-$\Delta$” transform; then, the QN becomes a series-parallel graph shown in Fig. S3(b), with three new TMSVSs of ratio negativity $\xi$ shared between $S$ and $R_{1}$, $R_{1}$ and $T$, and $S$ and $T$, respectively. Here, $\xi$ satisfies (by the star-mesh transform):
$\mathrm{seri}(\chi,\chi)=\mathrm{para}(\xi,\xi^{2}),$ (II.14)
@@
$y(x)=\sqrt{\frac{1+2\sqrt{1+3x^{-4}}\cos{\frac{\theta+\pi}{3}}}{3}},$ (II.16)
and $\theta=\arccos{\left[2^{-1}(-9x^{-4}+25)(1+3x^{-4})^{-\frac{3}{2}}\right]}$. Specifically, let $W=\chi^{4}$ and $Y=\xi^{2}$. Then, Eq. (II.15) is equivalent to
$Y^{3}-Y^{2}-W^{-1}Y+1=0,$ (II.17)
where $0\leqslant Y\leqslant W\leqslant 1$. The Shengjin’s formula [94] implies that the roots of this cubic polynomial are
$\left\{\begin{array}[]{l}Y_{1}={\left[1-2\sqrt{1+3W^{-1}}\cos{\frac{\theta}{3}}\right]}/{3},\\ Y_{2,3}={\left[1+2\sqrt{1+3W^{-1}}\left(\cos{\frac{\theta\mp\pi}{3}}\right)\right]}/{3},\\ \end{array}\right.$ (II.18)
where $\theta=\arccos\left[2^{-1}(-9W^{-1}+25)(1+3W^{-1})^{-\frac{3}{2}}\right]$. Only the third root $Y_{3}$ meets the numerical requirements $0\leqslant Y\leqslant W\leqslant 1$. Therefore, we have $\xi=\sqrt{Y_{3}}=y(\chi)$.
Solving Fig. S3(b), the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ is the ratio negativity of the final TMSVS between node $S$ and $T$ [Fig. S3(c)], which is given by $\mathrm{X}_{\text{SC}}=b(\chi,y(\chi))$ with
@@
Figure S4: Kelvin bridge. S4(a) The original QN. S4(b) The QN after the “$\Delta$-Y” transform. S4(c) The QN after applying series rules to $R_{1}$, $R_{2}$ and $R_{3}$. S4(d) The final QN.
B Kelvin bridge

### v2

@@ Methods
Lemma 1 provides a sufficient condition to convert two TMSVSs $|\psi^{r_{1}}\rangle$ and $|\psi^{r_{2}}\rangle$ into another one with squeezing parameter $r_{1}^{\prime}>r_{1}$ for the case of $r_{2}^{\prime}=0$. To be exact, the maximum value $r$ of $r_{1}^{\prime}$ that satisfies Eq. (14) is given by $r=\sinh^{\mbox{-}1}(\sinh{r_{1}}\cosh{r_{2}})>r_{1}$ and $r_{2}^{\prime}=0$ with $G=\eta^{-1}=(1-\tanh r_{2})^{-1}$, implying that entanglement is concentrated. Thus, consider a $(2+2)$-mode pure Gaussian state $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ with decreasingly ordered squeezing vectors $\vec{r}^{\downarrow}=(r_{1},r_{2})$. This state can be deterministically converted into a single TMSVS $|\psi^{r}\rangle$ by non-Gaussian LOCC, where
$\sinh{r}=\sinh{r_{1}}\cosh{r_{2}}.$ (15)
That is, according to Ref. [51], the bipartite state $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$ can be converted into $|\psi^{r}\rangle$ ($r_{1}<r$) via deterministic LOCC when the reduced state of $|\psi^{r}\rangle$ can be transformed into the reduced state of $|\psi^{r_{1}}\rangle$ through a pure loss channel with intensity transmittance $\eta=1-\tanh^{2}r_{2}$ ($r_{2}\leq r_{1}$).
Equation (15) is the building block for scaling the concentration protocol. Let $|\Psi\rangle$ be a pure Gaussian state and, w.l.o.g., write $|\Psi\rangle=\mathop{\bigotimes}\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$ $(r_{1}\geq r_{2}\geq\cdots\geq r_{K}>0)$ with squeezing vector $\vec{r}^{\downarrow}=\left(r_{1},r_{2},\dots,r_{K}\right)$ by Ref. [16]. By iteratively applying Eq. (15) K1K-1 times on $|\Psi\rangle$ (Fig. 6), the pure Gaussian state $|\Psi\rangle$ is deterministically concentrated into a single TMSVS. Only LOCC is used throughout the entire process, giving rise to the following theorem:
@@
$\sinh{r}=\sinh{r_{1}}\prod_{k=2}^{K}\cosh{r_{k}}.$ (16)
@@
Figure 6: Continuous-variable entanglement concentration. Consider $K$ parallel TMSVSs written as the tensor product: $\bigotimes\nolimits_{k=1}^{K}|\psi^{r_{k}}\rangle$. Step $a_{1}$: The TMSVS $|\psi^{r_{1}^{\prime}}\rangle$ is obtained by executing the TMSVS entanglement concentration protocol on $|\psi^{r_{1}}\rangle\otimes|\psi^{r_{2}}\rangle$; Step $a_{2}$: Performing the scheme again on $|\psi^{r_{1}^{\prime}}\rangle\otimes|\psi^{r_{3}}\rangle$ yields the TMSVS $|\psi^{r_{2}^{\prime}}\rangle$; and so on. This eventually results in a single TMSVS $|\psi^{r}\rangle$, where the squeezing parameter $r$ satisfies Eq. (16).
We now illustrate the optimal concentration order. In general, depending on which two modes are first concentrated, the prefactor in Eq. (16) could be $\sinh{r_{k}}$ for any $k$ other than $\sinh{r_{1}}$. However, the optimal entanglement of $|\psi^{r}\rangle$ is obtained by first concentrating the mode of the highest entanglement (largest squeezing parameter, $r_{1}$):
@@
The squeezing parameter $r$ of the final TMSVS $|\psi^{r}\rangle$ is maximized if and only if the TMSVS with the largest squeezing parameter ($\max\limits_{1\leq k\leq K}r_{k}$) is among the first two TMSVSs to be concentrated. The corresponding optimal $r$ is given by $\sinh r=(\tanh{\max\limits_{1\leq k\leq K}r_{k})}\left(\prod_{k=1}^{K}\cosh r_{k}\right)$.
@@
where $r^{\prime}=\text{arcsinh}(\sinh r_{2}\cosh r_{3})$. Since $r_{1}\geq r_{k}$ for $k=2,3$, we have $\sin r_{1}\cosh r_{2}\cosh r_{3}\geq\cosh r_{1}\sinh r^{\prime}$. Furthermore, $\sin r_{1}\cosh r_{2}\cosh r_{3}\geq\sinh r_{1}\cosh r^{\prime}$ holds, due to
@@
Case 2 ($K>3$). Assume that this theorem holds true for $K=n\ (n\geq 3)$. Then, for $K=n+1$, $|\psi^{r_{1}}\rangle$ must either be the first to concentrate (since it is the largest among the first $n$ TMSVSs to be concentrated), or the last to concentrate. In the latter case, $|\psi^{r_{2}}\rangle$ (with the second largest $r_{2}$), which is now the largest among the first $n$ TMSVSs, must be the first to be concentrated. This gives rise to
@@
where $r^{\prime\prime}=\text{arcsinh}\left(\sinh r_{2}\prod_{k=3}^{n+1}\cosh r_{k}\right)$. Because $r_{1}\geq r_{2}$ and
@@
$\leq$ $\prod_{k=2}^{n+1}\cosh r_{k},$
we have $\sinh r_{1}\prod_{k=2}^{n+1}\cosh r_{k}\geq\cosh r_{1}\sinh r^{\prime\prime}$ and $\sinh r_{1}\prod_{k=2}^{n+1}\cosh r_{k}\geq\sinh r_{1}\cosh r^{\prime\prime}$, respectively. It is therefore optimal to concentrate $|\psi^{r_{1}}\rangle$ first. ∎
To express Eq. (16) in terms of entanglement measure, let $\chi_{k}$ and $\chi$ be the ratio negativity of input TMSVS $|\psi^{r_{k}}\rangle$ and the output TMSVS $|\psi^{r}\rangle$, respectively. We have:
$\frac{\chi^{2}}{1-\chi^{2}}=\frac{\max\limits_{1\leq k\leq K}\chi_{k}^{2}}{\prod^{K}_{k=1}\left(1-\chi_{k}^{2}\right)},$ (17)
@@
$\text{para}(\chi_{1},\chi_{2},...):=\chi=\frac{\max\limits_{1\leq k\leq K}\chi_{k}}{\sqrt{\max\limits_{1\leq k\leq K}\chi_{k}^{2}+\prod_{k=1}^{K}(1-\chi_{k}^{2})}}.$
.3 Negativity Percolation in Non-Series-Parallel QNs
For non-series-parallel QNs, the negativity percolation theory (NegPT) based on the ratio negativity is defined using approximate star-mesh transforms (more details are provided in Ref. [66]). A star-mesh transform [83] can establish a local equivalence of the connectivity between an $(s+1)$-node star graph and an $s$-node complete graph, based only on series and parallel rules. Note that the star-mesh transform is generally irreversible. Specifically, while a star graph with arbitrary weights can most likely be transformed into a complete graph with matching weights, the inverse transformation from a complete graph with arbitrary weights back to a star graph is usually not possible. An exception to this is when $s=3$, in which case the star-mesh transform is reversible and is known as the “Y-$\Delta$$\Delta$-Y” transforms.
@@
Figure 7: Wheatstone bridge. (a) The original QN. (c) The QN after the “Y-$\Delta$” transform. (c) The final QN.
From Eqs. (10) and (.2), we have
@@
The simplest non-series-parallel network is a 4-node network: a Wheatstone bridge, as shown in Fig. 7(a). Let each link in the QN be a TMSVS with identical ratio negativity $\chi$. We apply the star-mesh transform to $R_{2}$, namely the “Y-$\Delta$” transform; then, the QN becomes a series-parallel graph shown in Fig. 7(b), with three new TMSVSs of ratio negativity $\xi$ shared between $S$ and $R_{1}$, $R_{1}$ and $T$, and $S$ and $T$, respectively. Here, $\xi$ satisfies (by the star-mesh transform):
$\mathrm{seri}(\chi,\chi)=\mathrm{para}(\xi,\xi^{2}),$ (19)
@@
$y(x)=\sqrt{\frac{1+2\sqrt{1+3x^{-4}}\cos{\frac{\theta+\pi}{3}}}{3}},$ (21)
and $\theta=\arccos{\left[2^{-1}(-9x^{-4}+25)(1+3x^{-4})^{-\frac{3}{2}}\right]}$. Specifically, let $W=\chi^{4}$ and $Y=\xi^{2}$. Then, Eq. (20) is equivalent to
$Y^{3}-Y^{2}-W^{-1}Y+1=0,$ (22)
where $0\leq Y\leq W\leq 1$. The Shengjin’s formula [36] implies that the roots of this cubic polynomial are
[…]
where $\theta=\arccos\left[2^{-1}(-9W^{-1}+25)(1+3W^{-1})^{-\frac{3}{2}}\right]$. Only the third root $Y_{3}$ meets the numerical requirements $0\leq Y\leq W\leq 1$. Therefore, we have $\xi=\sqrt{Y_{3}}=y(\chi)$.
Solving Fig. 7(b), the sponge-crossing ratio negativity $\mathrm{X}_{\text{SC}}$ is the ratio negativity of the final TMSVS between node $S$ and $T$ [Fig. 7(c)], which is given by $\mathrm{X}_{\text{SC}}=b(\chi,y(\chi))$ with
@@
Figure 8: Kelvin bridge. (a) The original QN. (b) The QN after the $\Delta$-Y” transform. (c) The QN after applying series rules to $R_{1}$, $R_{2}$ and $R_{3}$. (d) The final QN.